Sipacore Hospital is a division of Superior Healthcare that is organized as an investment center. In the past year, the hospital reported an after-tax income of $3,600,000. Total interest expense was $1,700,000, and the hospital’s tax rate was 35 percent. Hospital assets totaled $35,300,000, and noninterest-bearing current liabilities were $11,100,000. Superior has established a required rate of return equal to 17 percent of invested capital.

Calculate the residual income/EVA of Sipacore Hospital. (Enter negative answers preceding either - sign, e.g. -45 or in parentheses, e.g. (45).)
Answer #1

Answer:

Net operating income = 3,600,000 + 1,700,000 * (1 - 35%)

= 5,300,000 * (1 - 0.35)

= 3,445,000

Net operating assets = 35,300,000 - 11,100,000

= 24,200,000

Residual income = Net operating income - (Net operating assets * Required rate of return)

= 3,445,000 - (24,200,000 * 17%)

= 3,445,000 - 4,114,000

= (669,000) or - 669,000